# Study of hybrid stars with nonstrange quark matter cores

### v2

@@ Abstract
In this work, under the hypothesis that quark matter may not be strange [Phys. Rev. Lett. 120, 222001 (2018)], we adopt a modification of the coupling constant of the four-quark scalar interaction $G\rightarrow G_{1}+G_{2}\langle\bar{\psi}\psi\rangle$ in the 2-flavor Nambu-Jona-Lasinio model to study nonstrange hybrid stars. According to lattice QCD simulation results of the critical temperature at zero chemical potential, $G_{1}$ and $G_{2}$ are constrained as $G_{1}\in(1.935,1.972)$ GeV-2, and $G_{2}\in(-1.582,-0.743)$ GeV-5, respectively. To obtain hybrid equation of states, the Maxwell construction is used to describe the first-order confinement-deconfinement phase transition in hybrid stars. With recent measurements on neutron star mass, radius, and tidal deformability, the hybrid equation of states are constrained. The result suggests that pure nonstrange quark matter cores can exist in hybrid stars, possessing 0.014-0.026 solar mass with the bag constant $B^{1/4}$ in a range of 148-161 MeV. It is argued that the binary neutron stars in GW170817 should be hadron stars.
Key-words: hybrid star, Nambu-Jona-Lasinio model, equation of state, Maxwell construction
PACS Numbers: 12.38.Lg, 25.75.Nq, 21.65.Mn
pacs:
12.38.Mh, 12.39.-x, 25.75.Nq
I INTRODUCTION
@@
The binary neutron star (BNS) merger GW170817 opens a new era of multi-messenger astronomy Abbott and et al. (2017a, b); Margalit and Metzger (2017); Bauswein et al. (2017); Shibata et al. (2017); Annala et al. (2018); Fattoyev et al. (2018); Paschalidis et al. (2018); Zhou et al. (2018); Ruiz et al. (2018); Radice et al. (2018); Rezzolla et al. (2018); Nandi and Char (2018); Zhu et al. (2018); Ai et al. (2018); Ma et al. (2019); Zhang (2020); Li et al. (2020); Miao et al. (2021); Tan et al. (2022); Zou and Huang (2022); Li et al. (2022a); Zhang et al. (2023). More and more astronomical observations on neutron stars arise, facilitating the study of neutron star structure and equation of state (EOS). As natural laboratories to investigate the dense strongly interacting matter, neutron stars have been attracting much attention in astrophysics and theoretical physics. In general, the characteristic temperature of neutron stars can be well described by zero temperature approximation, due to their excessively high energy density in the interior, thus the quantum chromodynamics (QCD) needs to be employed to study the EOS in neutron stars. It is believed that the density in the core of neutron stars could reach 5-10 $\rho_{0}$, where $\rho_{0}=0.16$ fm-3 is the nuclear saturation density Lattimer and Prakash (2004); Lattimer (2021). As a result, the hadron-quark phase transition is very likely to happen and the deconfined quark matter will appear. In this case, neutron stars are essentially hybrid stars. However, it is difficult to give a unified description of the hadronic matter, quark matter and the hadron-quark phase transition with a single theoretical framework. Thus the hadronic matter and quark matter in hybrid stars are separately described with different EOSs at present, and a certain construction scheme needs to be employed to combine them to get a complete EOS.
@@
As we know, the results of different effective models can be quantitatively or even qualitatively different. Even for the same model, if different modifications are taken into account, the results can also be different. For example, Fig. 10 of Ref. Özel and Freire (2016) shows that the EOSs given by different effective models are different from each other, and the corresponding mass-radius ($M-R$) relations of neutron stars are also different. Thus there is not a definite answer to the EOS of dense strongly interacting matter at zero temperature at present.
@@
To describe the hadronic matter in hybrid stars, the Hartree-Fock-Bogoliubov (HFB) mass model HFB-21 with unconventional Skyrme forces called BSK21 Goriely et al. (2010); Chamel et al. (2011); Fantina, A. F. et al. (2013); Giliberti and Cambiotti (2022) is a good approach, and the corresponding hadronic EOS obtained in this framework is consistent with current constraints on neutron star masses, radii, and the tidal deformability from GW170817 Fantina, A. F. et al. (2022). However, to describe the quark matter in hybrid stars, the lattice QCD is confronted with difficulties at low-temperature and high-density regions because of the “sign problem”, thus we need to use effective models, such as the Nambu-Jona-Lasinio (NJL) model Klevansky (1992); Buballa (2005); Li et al. (2019a); Liu et al. (2021); Zhang et al. (2022); Huang and Zhuang (2023), which manifests the spontaneous breaking of chiral symmetry.
@@
In the framework of the NJL-type model, many studies focused on hybrid stars Ayriyan et al. (2021); Blaschke et al. (2020); Alvarez-Castillo et al. (2016); Pfaff et al. (2022), aiming to explain the observed two-solar-mass (2 $M_{\odot}$) compact stars. In Refs. Ayriyan et al. (2021); Blaschke et al. (2020); Alvarez-Castillo et al. (2016), the 2-flavor NJL-type model was used to consider the scalar quark-antiquark interaction, anti-triplet scalar diquark interactions and vector quark-antiquark interactions Ayriyan et al. (2021), a chemical potential dependence of the vector mean-field coupling $\eta(\mu)$ and a chemical potential-dependent bag constant $B(\mu)$ Blaschke et al. (2020), multiquark (4- and 8-quark) interactions Alvarez-Castillo et al. (2016), respectively. In Ref. Pfaff et al. (2022), the 3-flavor SU(3) NJL model was adopted with the four-quark scalar, vector-isoscalar and vector-isovector interactions as well as the ’t Hooft interaction. Different from the above studies, here we adopt a modification of the coupling constant of four-quark scalar interactions as $G\rightarrow G_{1}+G_{2}\langle\bar{\psi}\psi\rangle$ in this work, which can be regarded as a representation of an effective gluon propagator (See Sec. II for specific analysis).
@@
As for the hadron-quark phase transition, the most widely used approach is the Maxwell construction Endo et al. (2006); Hempel et al. (2009); Yasutake et al. (2009), assuming that the first-order phase transition occurs Glendenning (1992); Bhattacharyya et al. (2010) and stable quark matter cores exist in hybrid stars. However, many studies showed that hybrid stars are unstable against oscillations in this case, because star masses decrease with the increase of the central density, thus quark matter cores may not exist in neutron-star interiors Özel (2006); Hoyos et al. (2016); Qin et al. (2023). In Ref. Özel (2006) and Ref. Hoyos et al. (2016), the theoretical modeling of bursting neutron-star spectra and top-down holographic model for strongly interacting quark matter were employed, respectively, to demonstrate that the 2 $M_{\odot}$ neutron star has ruled out soft EOSs of neutron-star matter, and no quark matter exists in massive neutron stars. Recently, it has been argued that as the density increases, the boundaries of hadrons disappear gradually and the corresponding phase transition is a crossover Baym et al. (2018). According to this assumption, the three-window modeling Masuda et al. (2013a, b) in the crossover region was proposed. Many studies has constructed hybrid EOSs in this scheme, and the corresponding maximum masses of hybrid stars are compatible with 2 $M_{\odot}$ojo et al. (2015); Li et al. (2017, 2018a, 2018b, 2022b); Qin et al. (2023).
@@
In addition to theoretical studies of hadron-quark phase transitions and hybrid EOSs, astronomical observations of neutron star masses, radii, and tidal deformability have also placed constraints on numerous EOSs. Some massive neutron star observations such as PSR J0348+0432 Antoniadis et al. (2013) and PSR J0740+6620 Cromartie et al. (2020) require EOSs should not be too soft, but the tidal deformability constrained in BNS merger event GW170817 indicates the EOSs should not be too stiff Abbott and et al. (2017a, 2018). Recently, the joint $M-R$ observations of neutron stars from NASA’s Neutron Star Interior Composition Explorer (NICER) missions have also imposed some constraints on these EOSs Riley et al. (2019); Miller et al. (2019a); Riley et al. (2021); Miller et al. (2021). In Refs. Bauswein et al. (2019); Miao et al. (2020), the authors claim that the gravitational-wave (GW) emission of GW170817 supports a first-order hadron-quark phase transition at supranuclear densities.
@@
In this work, inspired by a recent work that the quark matter may not be strange Holdom et al. (2018), we will study nonstrange hybrid EOSs and hybrid stars with the Maxwell construction. The hadronic EOS and quark EOS are described by the HFB mass model BSK21 and a modified 2-flavor NJL model, respectively. The parameter space of $G_{1}$ and $G_{2}$ in the modified NJL model will be fixed according to the lattice results at zero chemical potential Laermann and Philipsen (2003). With recent measurements on neutron star mass, radii, and tidal deformability, the hybrid EOSs will be constrained to get the parameter space of the bag constant $B^{1/4}$. To ensure that hybrid stars are stable against oscillations, maximum masses of hybrid stars and the masses of their quark matter cores are determined.
@@
It is known that the Bayesian analysis is a good approach to constrain the EOSs. Researchers have obtained important information about the EOS of QCD in this way Ayriyan et al. (2021); Blaschke et al. (2020); Alvarez-Castillo et al. (2016); Pfaff et al. (2022); Alvarez-Castillo et al. (2020); Ayriyan et al. (2019, 2015); Miller et al. (2019b). However, considering that the lagrangian of the NJL model is convenient for numerical calculation, in this work we have performed calculations focusing on the EOS to get the corresponding hybrid star $M-R$ relations and tidal deformability, and then compared them with the relevant neutron star astronomical observations. The model parameters and properties of hybrid stars are constrained as well.
@@
This paper is organized as follows. In Sec.II, the modified NJL model for nonstrange quark matter is briefly introduced, and the EOSs of quark matter are derived. In Sec.III, the Maxwell construction is used to get hybrid EOSs. With recent astronomical observations of neutron star mass, radius, and tidal deformability, we constrain the hybrid EOSs. For comparison, the $M-R$ relations and tidal deformability results of hybrid stars from six representative hybrid EOSs are presented. Finally, a brief summary is given in Sec.IV.
II EOS of nonstrange quark matter
@@
As an effective model to describe cold dense quark matter, the NJL model Klevansky (1992); Buballa (2005) is widely used in the study of hybrid stars and quark stars. The general form of the 2-flavor NJL model Lagrangian is:
@@
where $m$ is the current quark mass (because of an exact isospin symmetry between $u$ and $d$ quarks adopted in this work, $m_{\rm u}=m_{\rm d}=m$), and $G$ represents the four-fermion coupling constant. The term $G[(\bar{\psi}\psi)^{2}+(\bar{\psi}i\gamma^{5}\tau\psi)^{2}]$ describes interactions in scalar-isoscalar and pseudoscalar-isovector channels.
Then the mean-field thermodynamic potential is
$\Omega(T,\{\mu_{\rm f}\},\{\langle\bar{\psi}\psi\rangle_{\rm f}\})=$ $\sum_{\rm f=u,d}(\Omega_{\rm M,f}(T,\mu_{\rm f})+2G\langle\bar{\psi}\psi\rangle_{\rm f}^{2})$ (2)
$+const,$
where $\langle\bar{\psi}\psi\rangle_{\rm f}$ is the quark condensate of flavor $f$, and $\Omega_{\rm M,f}$ denotes the contribution of a gas of quasiparticles of flavor $f$,
$\Omega_{\rm M,f}=$ $-2N_{\rm c}\int\frac{{\rm d}^{3}p}{(2\pi)^{3}}\{T\,{\rm ln}(1+{\rm exp}(-\frac{1}{T}(E_{p,{\rm f}}-\mu_{\rm f})))$ (3)
$+T\,{\rm ln}(1+{\rm exp}(-\frac{1}{T}(E_{p,{\rm f}}+\mu_{\rm f})))+E_{p,{\rm f}}\}.$
The effective mass of the constituent quark of flavor $f$ is now given by
$M_{\rm f}$ $=$ $m_{\rm f}-4G\langle\bar{\psi}\psi\rangle_{\rm f}.$ (4)
In a thermodynamically consistent treatment, the quark condensate $\langle\bar{\psi}\psi\rangle_{\rm f}$ and the particle number density $\rho_{\rm f}$ of flavor $f$ can be derived from $\Omega$ as
$\langle\bar{\psi}\psi\rangle_{\rm f}$ $=$ $\frac{\partial\Omega}{\partial m_{\rm f}}$
$=$ $-2N_{\rm c}\int\frac{{\rm d}^{3}p}{(2\pi)^{3}}\frac{M_{\rm f}}{E_{p,{\rm f}}}[1-n_{p,{\rm f}}(T,\mu_{\rm f})-\overline{n}_{p,{\rm f}}(T,\mu_{\rm f})],$
$\rho_{\rm f}$ $=$ $-\frac{\partial\Omega}{\partial\mu_{\rm f}}$ (6)
$=$ $2N_{\rm c}\int\frac{{\rm d}^{3}p}{(2\pi)^{3}}(n_{p,{\rm f}}(T,\mu_{\rm f})-\overline{n}_{p,{\rm f}}(T,\mu_{\rm f})),$
where $E_{p,{\rm f}}=\sqrt{\overrightarrow{p}^{2}+M_{\rm f}^{2}}$ is the quark on-shell energy of flavor $f$, and $n_{p,{\rm f}}(T,\mu_{\rm f})$, $\overline{n}_{p,{\rm f}}(T,\mu_{\rm f})$ are the Fermi occupation numbers of quarks and antiquarks of flavor $f$, respectively, which are defined as
$n_{p,{\rm f}}(T,\mu_{\rm f})$ $=$ $[{\rm exp}^{(E_{p,{\rm f}}-\mu_{\rm f})/T}+1]^{-1},$ (7)
$\overline{n}_{p,{\rm f}}(T,\mu_{\rm f})$ $=$ $[{\rm exp}^{(E_{p,{\rm f}}+\mu_{\rm f})/T}+1]^{-1}.$ (8)
Because the NJL model cannot be renormalized, the proper-time regularization is adopted in the following calculations. In addition, we need to fix the parameter set ($\Lambda_{\rm UV},G$) to fit experimental data ($f_{\pi}=92$ MeV, $M_{\pi}=135$ MeV) at zero temperature and chemical potential. According to the Review of Particle Physics Particle Data Group et al. (2022), the current quark mass $m$ is chosen as 3.5 MeV in this work. Then the ultraviolet cutoff $\Lambda_{\rm UV}$ and coupling constant $G$ are fixed to be 1324 MeV and 2.005 GeV-2, respectively. This procedure is similar to Ref. Klevansky (1992).
@@
Based on our current knowledge of strong interactions, the coupling constant $G$ in the NJL model can be regarded as a representation of an effective gluon propagator. In light of QCD theory, the quark and gluon propagators should satisfy their respective Dyson-Schwinger (DS) equations, and these two equations are coupled with each other. It is demonstrated that quark propagators in the Nambu phase and Wigner phase are very different from each other Cui et al. (2018); Xu et al. (2018); Li et al. (2019b), so it can be inferred that the corresponding gluon propagators in these two phases are also different Hong-Shi and Wei-Min (2006). However, in the normal NJL model, $G$ is simplified as a constant, remaining the same in these two phases. In addition, according to simulations of lattice QCD, the gluon propagator should vary with temperature, although its dependence on the chemical potential is still uncertain. In the normal NJL model, as a representation of an effective gluon propagator, the coupling constant $G$ is “static”, and thus cannot fulfill the requirement of lattice QCD.
@@
In the QCD sum rule approach Reinders et al. (1985), it is argued that the full Green function can be divided into two parts: the perturbative part and nonperturbative part. The condensates can be expressed as various moments of nonperturbative Green function. As a result, the most general form of the “nonperturbative” gluon propagator is
@@
where $\langle G^{\mu\nu}G_{\mu\nu}\rangle$ refers to the gluon condensate, $c_{1}$ and $c_{2}$ are coefficients which can be calculated in the QCD sum rule approach Steele (1989); Pascual and Tarrach (1984), and the ellipsis represents the contributions from other condensates, such as the mixed quark-gluon condensate. Among these condensates, the quark condensate possesses the lowest dimension, and a nonzero value of it, in the chiral limit, precisely signifies the dynamical chiral symmetry breaking. Therefore, it plays the most important role in the QCD sum rule approach. In this work, we will deal with its contribution separately, and the contribution of other condensates is simplified into the perturbative part of the gluon propagator. In the normal NJL model, it is equivalent to a modification of the coupling constant $G$ in the following way Jiang et al. (2012); Cui et al. (2013, 2014); Shi et al. (2016); Wang et al. (2016); Fan et al. (2017, 2019); Li et al. (2018a),
@@
Now the coupling strength $G$ will depend on both $u$ and $d$ quark condensates via this modification, where $G_{2}$ refers to the weight factor of the influence of the quark propagator on the gluon propagator.
@@
Although the lattice QCD is confronted with the “sign problem” at finite chemical potentials, the simulating results at zero chemical potential can still help us determine the values of $G_{1}$ and $G_{2}$. According to the simulations of lattice QCD, the chiral phase transition at zero chemical potential is a crossover, and the corresponding pseudo-critical point is located at $T_{\rm pc}=173\pm 8$ MeV in the 2-flavor caseLaermann and Philipsen (2003). Different from the meaning of the so-called “critical point” in the case of first-order phase transition, the “pseudo-critical point” here refers to the condition that the crossover occurs, and its position can be identified by the peak of susceptibilities, such as the chiral susceptibility $\chi_{s}$ in Ref. Laermann and Philipsen (2003), which is defined as $\chi_{s}=-\partial\langle\bar{\psi}\psi\rangle/\partial m$ Du et al. (2013).

### v3

@@ Abstract
In this work, under the hypothesis that quark matter may not be strange (Holdom et al., 2018), we adopt a modification of the coupling constant of the four-quark scalar interaction $G\rightarrow G_{1}+G_{2}\langle\bar{\psi}\psi\rangle$ in the 2-flavor Nambu-Jona-Lasinio (NJL) model to study nonstrange hybrid stars, where $G_{1}$ and $G_{2}$ are two parameters constrained by using the lattice QCD simulation results at the critical temperature and zero chemical potential. The Maxwell construction is used to describe the first-order confinement-deconfinement phase transition in hybrid stars. With recent measurements on neutron star mass, radius, and tidal deformability, the hybrid equation of states are constrained. It is found that pure nonstrange quark matter cores can exist in hybrid stars, possessing $0.026-0.04$ solar mass. The maximum hybrid star mass in the framework of the modified NJL model is about 0.1 solar mass lighter than that in the conventional 2-flavor NJL model. It is argued that the binary neutron stars in GW170817 should be hadron stars.
thanks: E-mail: licm@zzu.edu.cnthanks: E-mail: zhaoyapeng2013@hotmail.com
1 Introduction
@@
The binary neutron star (BNS) merger GW170817 opened a new era of multi-messenger astronomy  (Abbott & et al., 2017a, b; Margalit & Metzger, 2017; Bauswein et al., 2017; Shibata et al., 2017; Annala et al., 2018; Fattoyev et al., 2018; Paschalidis et al., 2018; Zhou et al., 2018; Ruiz et al., 2018; Radice et al., 2018; Rezzolla et al., 2018; Nandi & Char, 2018; Zhu et al., 2018; Ai et al., 2018; Ma et al., 2019; Zhang, 2020; Li et al., 2020; Miao et al., 2021; Tan et al., 2022; Zou & Huang, 2022; Li et al., 2022b; Zhang et al., 2023). More and more astronomical observations on neutron stars arise, facilitating the study of neutron star structure and equation of state (EOS). As natural laboratories to investigate the dense strongly interacting matter, neutron stars have been attracting much attention in astrophysics and theoretical physics. In general, the characteristic temperature of neutron stars can be well described by zero temperature approximation, due to their excessively high energy density in the interior, thus the quantum chromodynamics (QCD) needs to be employed to study the EOS in neutron stars. It is believed that the density in the core of neutron stars could reach 5-10 $\rho_{0}$, where $\rho_{0}=0.16$ fm-3 is the nuclear saturation density (Lattimer & Prakash, 2004; Lattimer, 2021). As a result, the hadron-quark phase transition is very likely to happen and the deconfined quark matter will appear. In this case, neutron stars are essentially hybrid stars. However, it is difficult to give a unified description of the hadronic matter, quark matter and the hadron-quark phase transition with a single theoretical framework. Thus the hadronic matter and quark matter in hybrid stars are separately described with different EOSs at present, and a certain construction scheme needs to be employed to combine them to get a complete EOS.
@@
As we know, the results of different effective models can be quantitatively or even qualitatively different. Even for the same model, if different modifications are taken into account, the results can also be different. For example, Fig. 10 of Özel & Freire (2016) shows that the EOSs given by different effective models are different from each other, and the corresponding mass-radius ($M-R$) relations of neutron stars are also different. Thus there is not a definite answer to the EOS of dense strongly interacting matter at zero temperature at present.
@@
To describe the hadronic matter in hybrid stars, the EOS developed by Akmal, Pandharipande & Ravenhall (APR) with $A18+\delta\nu+UIX^{\ast}$ interaction is employed in this work (Akmal et al., 1998), in which the Argonne $\nu_{18}$ two-nucleon interaction and boost corrections to the two-nucleon interaction as well as the three-nucleon interaction are taken into account. The hadronic matter in the context of the APR model is one kind of charge-neutral and beta-stable fluid whose pressure and baryon chemical potential are equilibrated. However, to describe the quark matter in hybrid stars, the lattice QCD is confronted with difficulties at low-temperature and high-density regions because of the “sign problem”, thus we need to use effective models, such as the Nambu-Jona-Lasinio (NJL) model  (Klevansky, 1992; Buballa, 2005; Li et al., 2019b; Liu et al., 2021; Zhang et al., 2022; Huang & Zhuang, 2023), which manifests the spontaneous breaking of chiral symmetry.
@@
In the framework of the NJL-type model, many studies focused on hybrid stars (Ayriyan et al., 2021; Blaschke et al., 2020; Alvarez-Castillo et al., 2016; Pfaff et al., 2022), aiming to explain the observed two-solar-mass (2 $M_{\odot}$) compact stars. In Ayriyan et al. (2021); Blaschke et al. (2020); Alvarez-Castillo et al. (2016), the 2-flavor NJL-type model was used to consider the scalar quark-antiquark interaction, anti-triplet scalar diquark interactions and vector quark-antiquark interactions  (Ayriyan et al., 2021), a chemical potential dependence of the vector mean-field coupling $\eta(\mu)$ and a chemical potential-dependent bag constant $B(\mu)$  (Blaschke et al., 2020), multiquark (4- and 8-quark) interactions  (Alvarez-Castillo et al., 2016), respectively. In Pfaff et al. (2022), the 3-flavor SU(3) NJL model was adopted with the four-quark scalar, vector-isoscalar and vector-isovector interactions as well as the ’t Hooft interaction. Different from the above studies, we adopt a modification of the coupling constant of four-quark scalar interactions as $G\rightarrow G_{1}+G_{2}\langle\bar{\psi}\psi\rangle$, which can be regarded as a representation of an effective gluon propagator (See Sec. 2 for specific analysis).
@@
As for the hadron-quark phase transition, the most widely used approach is the Maxwell construction  (Endo et al., 2006; Hempel et al., 2009; Yasutake et al., 2009), assuming that the first-order phase transition occurs (Glendenning, 1992; Bhattacharyya et al., 2010) and stable quark matter cores exist in hybrid stars. However, many studies showed that hybrid stars are unstable against oscillations in this case, because star masses decrease with the increase of the central density, thus quark matter cores may not exist in neutron-star interiors  (Özel, 2006; Hoyos et al., 2016; Qin et al., 2023). In Özel (2006) and Hoyos et al. (2016), the theoretical modeling of bursting neutron-star spectra and top-down holographic model for strongly interacting quark matter were employed, respectively, to demonstrate that the 2 $M_{\odot}$ neutron star has ruled out soft EOSs of neutron-star matter, and no quark matter exists in massive neutron stars. Recently, it has been argued that as the density increases, the boundaries of hadrons disappear gradually and the corresponding phase transition is a crossover  (Baym et al., 2018). According to this assumption, the three-window modeling (Masuda et al., 2013a, b) in the crossover region was proposed. Many studies has constructed hybrid EOSs in this scheme, and the corresponding maximum masses of hybrid stars are compatible with 2 $M_{\odot}$  (Kojo et al., 2015; Li et al., 2017, 2018b, 2018a, 2022a; Qin et al., 2023).
@@
In addition to theoretical studies of hadron-quark phase transitions and hybrid EOSs, astronomical observations of neutron star masses, radii, and tidal deformability have also placed constraints on numerous EOSs. Some massive neutron star observations such as PSR J0348+0432 (Antoniadis et al., 2013) and PSR J0740+6620 (Cromartie et al., 2020) require EOSs should not be too soft, but the tidal deformability constrained in BNS merger event GW170817 indicates the EOSs should not be too stiff (Abbott & et al., 2017a, 2018). Recently, the joint $M-R$ observations of neutron stars from NASA’s Neutron Star Interior Composition Explorer (NICER) missions have also imposed some constraints on these EOSs  (Riley et al., 2019; Miller et al., 2019b; Riley et al., 2021; Miller et al., 2021). In Bauswein et al. (2019); Miao et al. (2020), the authors claim that the gravitational-wave (GW) emission of GW170817 supports a first-order hadron-quark phase transition at supranuclear densities.
@@
In this work, inspired by a recent work that the quark matter may not be strange (Holdom et al., 2018), we will study nonstrange hybrid EOSs and hybrid stars with the Maxwell construction. The hadronic EOS and quark EOS are described by the APR model and a modified 2-flavor NJL model, respectively. The parameter space of $G_{1}$ and $G_{2}$ in the modified NJL model will be fixed according to the lattice results at zero chemical potential  (Laermann & Philipsen, 2003). With recent measurements on neutron star mass, radii, and tidal deformability, the hybrid EOSs will be constrained to get the parameter space of the current quark mass. To ensure that hybrid stars are stable against oscillations, maximum masses of hybrid stars and the masses of their quark matter cores are determined.
@@
It is known that the Bayesian analysis is a good approach to constrain the EOSs. Researchers have obtained important information about the EOS of QCD in this way (Ayriyan et al., 2021; Blaschke et al., 2020; Alvarez-Castillo et al., 2016; Pfaff et al., 2022; Alvarez-Castillo et al., 2020; Ayriyan et al., 2019, 2015; Miller et al., 2019a). However, considering that the lagrangian of the NJL model is convenient for numerical calculation, in this work we have performed calculations focusing on the EOS to get the corresponding hybrid star $M-R$ relations and tidal deformability, and then compared them with the relevant neutron star astronomical observations. The model parameters and properties of hybrid stars are constrained as well.
@@
This paper is organized as follows. In Sec.2, the modified NJL model for nonstrange quark matter is briefly introduced, and the EOSs of quark matter are derived. In Sec.3, the Maxwell construction is used to get hybrid EOSs. With recent astronomical observations of neutron star mass, radius, and tidal deformability, we constrain the hybrid EOSs. For comparison, the $M-R$ relations and tidal deformability results of hybrid stars from six representative hybrid EOSs are presented. Finally, a brief summary is given in Sec.4.
2 EOS of nonstrange quark matter
@@
As an effective model to describe cold dense quark matter, the NJL model (Klevansky, 1992; Buballa, 2005) is widely used in the study of hybrid stars and quark stars. In this work, we consider a modified version, in which the Lagrangian has the following form:
@@
where $m$ is the current quark mass (because of an exact isospin symmetry between $u$ and $d$ quarks adopted in this work, $m_{\rm u}=m_{\rm d}=m$). Different from the normal NJL model, we adopt $G_{1}+G_{2}\langle\bar{\psi}\psi\rangle$ to represent the four-fermion coupling strength, where $\langle\bar{\psi}\psi\rangle$ is the quark condensate. The term $(G_{1}+G_{2}\langle\bar{\psi}\psi\rangle)[(\bar{\psi}\psi)^{2}+(\bar{\psi}i\gamma^{5}\tau\psi)^{2}]$ describes interactions in scalar-isoscalar and pseudoscalar-isovector channels.
@@
In the following we will clarify the modification in detail. Based on our current knowledge of strong interactions, the coupling constant $G$ in the normal NJL model can be regarded as a representation of an effective gluon propagator. In light of QCD theory, the quark and gluon propagators should satisfy their respective Dyson-Schwinger (DS) equations, and these two equations are coupled with each other. It is demonstrated that quark propagators in the Nambu phase and Wigner phase are very different from each other  (Cui et al., 2018; Xu et al., 2018; Li et al., 2019a), so it can be inferred that the corresponding gluon propagators in these two phases are also different (Hong-Shi & Wei-Min, 2006). However, in the normal NJL model, $G$ is simplified as a constant, remaining the same in these two phases. In addition, according to simulations of lattice QCD, the gluon propagator should vary with temperature, although its dependence on the chemical potential is still uncertain. In the normal NJL model, as a representation of an effective gluon propagator, the coupling constant $G$ is “static”, and thus cannot fulfill the requirement of lattice QCD.
@@
In the QCD sum rule approach (Reinders et al., 1985), it is argued that the full Green function can be divided into two parts: the perturbative part and nonperturbative part. The condensates can be expressed as various moments of nonperturbative Green function. As a result, the most general form of the “nonperturbative” gluon propagator is
@@
where $\langle G^{\mu\nu}G_{\mu\nu}\rangle$ refers to the gluon condensate, $c_{1}$ and $c_{2}$ are coefficients which can be calculated in the QCD sum rule approach  (Steele, 1989; Pascual & Tarrach, 1984), and the ellipsis represents the contributions from other condensates, such as the mixed quark-gluon condensate. Among these condensates, the quark condensate possesses the lowest dimension, and a nonzero value of it, in the chiral limit, precisely signifies the dynamical chiral symmetry breaking. Therefore, it plays the most important role in the QCD sum rule approach. In this work, we will deal with its contribution separately, and the contribution of other condensates is simplified into the perturbative part of the gluon propagator. In the normal NJL model, it is equivalent to a modification of the coupling constant $G$ in the following way  (Jiang et al., 2012; Cui et al., 2013, 2014b; Shi et al., 2016; Wang et al., 2016; Fan et al., 2017, 2019; Li et al., 2018b),
@@
Now the coupling strength $G$ will depend on both $u$ and $d$ quark condensates via this modification, where $G_{2}$ refers to the weight factor of the influence of the quark propagator on the gluon propagator. It should be noted that following this modification, we need to correspondingly change the term $G\langle\bar{\psi}\psi\rangle^{2}$ in the mean-field thermodynamic potential of the normal NJL model to ensure the consistency in thermodynamics. Similar to Gorenstein & Yang (1995), the term $G\langle\bar{\psi}\psi\rangle^{2}$ is replaced by the integral over $\langle\bar{\psi}\psi\rangle$ of $2\langle\bar{\psi}\psi\rangle{\rm d}(G\langle\bar{\psi}\psi\rangle)/{\rm d}\langle\bar{\psi}\psi\rangle$, in which d…/d… denotes derivative, and the result is $G_{1}\langle\bar{\psi}\psi\rangle^{2}+4G_{2}\langle\bar{\psi}\psi\rangle^{3}/3$. In this way, the mean-field thermodynamic potential has the following form,
$\Omega(T,\{\mu\},\{\langle\bar{\psi}\psi\rangle\})=$ $\Omega_{\rm M}(T,\mu)+G_{1}\langle\bar{\psi}\psi\rangle^{2}+\frac{4}{3}G_{2}\langle\bar{\psi}\psi\rangle^{3}$ (4)
$+{\rm const},$
where $\Omega_{\rm M}$ denotes the contribution of a gas of quasiparticles,
$\Omega_{\rm M}=$ $-2N_{\rm c}N_{\rm f}\int\frac{{\rm d}^{3}p}{(2\pi)^{3}}\{T\,{\rm ln}(1+{\rm exp}(-\frac{1}{T}(E_{p}-\mu)))$ (5)
$+T\,{\rm ln}(1+{\rm exp}(-\frac{1}{T}(E_{p}+\mu)))+E_{p}\},$
where $E_{p}=\sqrt{\overrightarrow{p}^{2}+M^{2}}$ is the quark on-shell energy. Note that when $G={\rm const}$ (i.e., $G_{1}=G$ and $G_{2}=0$), our model reproduces the condensate term in the normal NJL model.
The effective mass of the constituent quark is now given by
$M$ $=$ $m-2(G_{1}+G_{2}\langle\bar{\psi}\psi\rangle)\langle\bar{\psi}\psi\rangle.$ (6)
The quark condensate $\langle\bar{\psi}\psi\rangle$ and the particle number density $\rho$ can be derived from $\Omega$ as
$\langle\bar{\psi}\psi\rangle$ $=$ $\frac{\partial\Omega}{\partial m}$
$=$ $-2N_{\rm c}N_{\rm f}\int\frac{{\rm d}^{3}p}{(2\pi)^{3}}\frac{M}{E_{p}}[1-n_{p}(T,\mu)-\overline{n}_{p}(T,\mu)],$
$\rho$ $=$ $-\frac{\partial\Omega}{\partial\mu}$ (8)
$=$ $2N_{\rm c}N_{\rm f}\int\frac{{\rm d}^{3}p}{(2\pi)^{3}}(n_{p}(T,\mu)-\overline{n}_{p}(T,\mu)),$
where $n_{p}(T,\mu)$ and $\overline{n}_{p}(T,\mu)$ are the Fermi occupation numbers of quarks and antiquarks, respectively, which are defined as
$n_{p}(T,\mu)$ $=$ $[{\rm exp}^{(E_{p}-\mu)/T}+1]^{-1},$ (9)
$\overline{n}_{p}(T,\mu)$ $=$ $[{\rm exp}^{(E_{p}+\mu)/T}+1]^{-1}.$ (10)
Because the NJL model cannot be renormalized, the proper-time regularization is adopted in the following calculations. In addition, we need to fix the parameter set ($\Lambda_{\rm UV}$, $G$) to fit experimental data ($f_{\pi}=92$ MeV, $M_{\pi}=135$ MeV) at zero temperature and chemical potential. The parameter fixing process is similar to that of Cui et al. (2014a).
@@
Although the lattice QCD is confronted with the “sign problem” at finite chemical potentials, the simulating results at zero chemical potential can still help us determine the values of $G_{1}$ and $G_{2}$. According to the simulations of lattice QCD, the chiral phase transition at zero chemical potential is a crossover, and the corresponding pseudo-critical point is located at $T_{\rm pc}=173\pm 8$ MeV in the 2-flavor case  (Laermann & Philipsen, 2003). Different from the meaning of the so-called “critical point” in the case of first-order phase transition, the “pseudo-critical point” here refers to the condition that the crossover occurs, and its position can be identified by the peak of susceptibilities, such as the chiral susceptibility $\chi_{s}$ in Laermann & Philipsen (2003), which is defined as $\chi_{s}=-\partial\langle\bar{\psi}\psi\rangle/\partial m$  (Du et al., 2013).